# Jian-hao Kanga, Zhun Goub and Nan-jing Huangc

### v1

@@ Abstract
This paper considers a newly delayed reinsurance and investment optimization problem incorporating random risk aversion, in which an insurer pursues maximization of the expected certainty equivalent of her/his terminal wealth and the cumulative delayed information of the wealth over a period. Specially, the insurer’s surplus dynamics are approximated using a drifted Brownian motion, while the financial market is described by the Black-Scholes model. Moreover, the performance-linked capital flow feature is incorporated and the wealth process is formulated via a stochastic delay differential equation (SDDE). By adopting a game-theoretic approach, a verification theorem with rigorous proofs is established to capture the equilibrium reinsurance and investment strategy along with the equilibrium value function. Furthermore, for the cases of exponential utility and power utility, analytical or semi-analytical equilibrium reinsurance and investment strategies together with their equilibrium value functions are obtained under mild conditions. Finally, several numerical experiments are conducted to analyze the behavioral characteristics of the freshly-derived equilibrium reinsurance and investment strategy.
@@ 1 Introduction
Risk management has always been a crucial subject for insurers. Reinsurance serves as an efficient mean for insurers to manage claim-related risk exposure and safeguard themselves against catastrophic loss scenarios. Within a reinsurance contract, an insurer transfers a specified portion of claim losses to a reinsurer and remits the reinsurance premium as consideration for the risk transfer. Conversely, the reinsurer has the obligation to reimburse a portion of the claims that the insurer has incurred. Generally, the magnitude of reinsurance premium rises in line with the proportion of ceded claims, which compels the insurer to strike a balance between the reinsurance premium and ceded claims in order to attain the optimally achieved state of the reinsurance arrangement. Moreover, nowadays, insurers are actively engaged in investment pursuits within financial markets. Thus, it is also highly relevant to formulate suitable investment strategies to boost returns to the fullest extent or/and cut down risks to the minimum when dealing with different investment tools.
A substantial number of studies have been conducted on formulating optimal reinsurance and investment strategies for insurers. In these studies, many scholars have dedicated themselves to expanding the previous research works to include a variety of optimality criteria and distinct foundational models for the insurer’s surplus dynamics and risky asset evolution. For example, Browne [13] studied the optimal portfolio allocation problem for an insurer, taking into account the criteria of utility maximization and minimization of ruin probability separately; Højgaard and Taksar [22] aimed at enabling an insurer to pursue maximization of the return function for the proportional reinsurance problem; Yang and Zhang [35] adopted a jump-diffusion model to capture an insurer’s surplus process and investigated the optimal investment issue. Many more studies and outcomes associated with the reinsurance and investment problems are available in the literature (see, for instance, Bai and Guo [4], Elliott and Siu [18], Li et al. [27], Wang et al. [33] and the references therein).
@@
One typical feature of the above researches is that these researches concentrated on the controlled state system in the absence of delay considerations. Specially, it was commonly supposed that the system’s future state has no connection with past states and is governed exclusively by the present situation. Nevertheless, in real-world systems, insurers often take past information into account when formulating their reinsurance and investment strategies, as past information can help insurers understand the cyclical patterns of the markets and enhance the robustness of decision-making. This feature is frequently termed as the memory feature or the delay feature. Chang et al. [15] used the HJB equation to study a consumption and investment optimization problem featuring bounded memory, where the return of the risky asset was associated with the historic performance. Then, Shen and Zeng [30] pioneered the incorporation of bounded wealth memory into a reinsurance and investment optimization problem subject to the mean-variance criterion, deriving the pre-committed strategy through the stochastic maximum principle. It was pointed out in [15, 30] that since the delayed optimal control problems are usually infinite-dimensional, a significant challenge emerges, namely, the non-existence of the explicit solutions. Therefore, some specific constraints on model parameters were given in [15, 30] to render the original problems finite-dimensional and solvable. Further research works and results related to delayed optimal control problems (including delayed optimal reinsurance/investment problems) are well-documented in the literature (see, for example, A et al. [1], A and Shao [2], A et al. [3], Bai et al. [5], Elsanosi et al. [19], Federico [20], Meng et al. [28], Yan and Wong [34], Yuan et al. [39] and the references therein).
@@
It is worthy of note that the majority of aforementioned studies assumed that investors/insurers had a constant risk aversion coefficient. In reality, investors/insurers cannot easily determine their exact level of risk aversion. This has motivated extensive research on risk aversion estimation through diverse methodologies such as employing data from the market on labor income [16], using data pertaining to implied and realized volatilities from the market [12] and examining survey forms of representative samples [14]. Recently, Desmettre and Steffensen [17] proposed that it is more reasonable to characterize the risk aversion coefficient of investors with a random variable and they formulated a new investment optimization problem in the Black-Scholes financial market by employing expected certainty equivalent, which is time inconsistent. They also provided a strict verification theorem to capture the equilibrium investment strategy, deriving (semi-)analytical strategies for both power and exponential utility specifications. More recently, based on the research of [17], Kang et al. [25] considered an optimal reinsurance and investment problem incorporating random risk aversion in the financial market characterized by the Heston stochastic volatility model. They also presented the pseudo HJB equation to characterize the equilibrium reinsurance and investment strategy, from which the semi-analytical equilibrium strategy was obtained under exponential utility.
@@
To the best knowledge of the authors, there is currently no literature that adopts the method of using expected certainty equivalent to study the delayed optimal reinsurance and investment problem under random risk aversion. Thus, this paper systematically addresses this gap by applying expected certainty equivalent to consider a new delayed optimal reinsurance and investment problem incorporating random risk aversion. In particular, under the diffusion approximation modeling of the insurer’s surplus process, the insurer can cede risk exposure to a reinsurer by employing proportional reinsurance strategy or seek growth through new business expansion and can also optimize asset allocations in a financial market following Black-Scholes dynamics. Furthermore, it is hypothesized that capital inflows and outflows affect the insurer’s present wealth, with the associated wealth dynamics formulated via an SDDE. Then, an approach based on expected certainty equivalent is employed to formulate the general delayed optimal reinsurance and investment problem considering random risk aversion that can be characterized as a combination of nonlinear functions of expectations.
@@
The above newly established problem exhibits time inconsistency because Bellman’s dynamic programming principle is inapplicable. In order to obtain the (semi-)analytical equilibrium reinsurance and investment strategy for such a problem, we identify the following challenges. Firstly, due to the introduction of delay, the methodologies in [17, 25] for providing the verification theorem are no longer directly applicable. In this paper, we draw on [19] to derive the Itô formula under delay scenario, and then adopt the methodologies from [9, 11, 17, 25], integrating them with the aforementioned Itô formula to provide a verification theorem in delay context. Secondly, as discussed earlier, the inclusion of delay leads to the problem being infinite-dimensional. To overcome this difficulty, we follow [15, 30] to impose specific assumptions on model parameters to transform the problem into a finite-dimensional framework, which facilitates our derivation of (semi-)analytical equilibrium reinsurance and investment strategies under exponential utility and power utility. Finally, we find that only a small number of studies, such as [5, 15], have proven the admissibility of their delayed optimal (or equilibrium) strategies and the conditions in their verification theorem. However, these proofs have certain limitations and cannot be directly applied to this paper. For example, substituting equations (56) and (57) into equation (12) in [15], we discover that the proof process of Lemma 2.2 is no longer applicable and thus $c^{*}(t)\geq 0$ in the proof of Theorem 4.2 cannot be easy to check. What is more, Bai et al. [5] did not consider the verification of the integrability of the value function (see the assumption $(A1)$ in this paper). To address this issue, we utilize the properties of solutions to stochastic differential equations (SDEs) and the assumption of model parameters to complete the corresponding proofs. After overcoming these challenges, we would like to point out that the newly derived equilibrium reinsurance and investment strategies extend some results in [1, 2, 17, 25] and also arrive at several novel and interesting conclusions. For instance, the expressions of the equilibrium reinsurance strategy and the equilibrium investment strategy under the power utility function demonstrate that even under the independence assumption between financial and insurance markets, both strategies are unexpectedly influenced by the model parameters from the reinsurance and financial market setups.
@@
The primary contributions of this paper can be outlined from four different aspects: (i) a newly delayed optimal reinsurance and investment model with random risk aversion is established, which fully generalizes the model in [17] and partially generalizes the model in [25], while also enriching the research on time-inconsistent problems under delay scenarios; (ii) a complete verification theorem is provided to capture the equilibrium reinsurance and investment strategy and the associated equilibrium value function for the aforementioned optimization problem; (iii) within exponential and power utility settings, analytical or semi-analytical equilibrium reinsurance and investment strategies together with the associated equilibrium value functions are obtained, followed by the proofs of their admissibility and the validity of the verification theorem’s conditions; (iv) computational experiments quantify parameter sensitivity in the equilibrium reinsurance and investment strategy, which yields novel and intriguing findings.
@@
The subsequent sections of this paper are structured as outlined below. Section 2 details the model framework and sets forth the fundamental assumptions. Section 3 poses the delayed optimal reinsurance and investment problem under random risk aversion and gives a verification theorem for capturing the equilibrium reinsurance and investment strategy along with the equilibrium value function. Section 4 obtains the (semi-)closed-form solutions to equilibrium reinsurance and investment strategies under exponential and power utilities separately. Section 5 presents numerical simulations to quantify how model parameters influence the equilibrium reinsurance and investment strategy. Section 6 offers a comprehensive summary of this study and Appendices include the detailed proofs.
@@ 2.2 Financial market
where $r>0$ is a constant denoting the risk-free interest rate. In addition, the price process $S(t)$ of the risky asset is governed by
$\mathrm{d}S(t)=S(t)[\mu\mathrm{d}t+\sigma\mathrm{d}W_{2}(t)],\quad S(0)=s>0,$
where positive constants $\mu$ and $\sigma$ characterize respectively the expected return rate and volatility of $S(t)$, and $W_{2}:=\{W_{2}(t)\}_{t\in[0,T]}$ is a standard $\mathbb{P}$-Brownian motion independent of $W_{1}$.
@@ 2.3 Wealth process with delay
$\mathrm{d}X^{u}(t)$ $=\frac{(1-\pi(t))X^{u}(t)}{S_{0}(t)}\mathrm{d}S_{0}(t)+\frac{\pi(t)X^{u}(t)}{S(t)}\mathrm{d}S(t)+\mathrm{d}R(t)$
$=[X^{u}(t)\left(r+\pi(t)(\mu-r)\right)+a\eta+a\eta_{2}q(t)]\mathrm{d}t+bq(t)\mathrm{d}W_{1}(t)+\pi(t)X^{u}(t)\sigma\mathrm{d}W_{2}(t)$
with the initial wealth $X^{u}(0)=x_{0}>0$.
@@
$\mathrm{d}X^{u}(t)=$ $[X^{u}(t)\left(r+\pi(t)(\mu-r)\right)+a\eta+a\eta_{2}q(t)]\mathrm{d}t+bq(t)\mathrm{d}W_{1}(t)+\pi(t)X^{u}(t)\sigma\mathrm{d}W_{2}(t)$
@@
$+bq(t)\mathrm{d}W_{1}(t)+\pi(t)X^{u}(t)\sigma\mathrm{d}W_{2}(t),$ (2.3)
@@
We suppose that the initial condition for $X^{u}(t)$ is its historical performance from $\tau-h$ to $\tau$ formulated as $X^{u}(t)=\psi(t-\tau)$ for $t\in[\tau-h,\tau]$, where $\psi>0$ is a continuous function on [h,0][-h,0]. If $\tau=0$, then we suppose that the insurer’s initial wealth is constant, namely, for any $t\in[-h,0]$, $X^{u}(t)=\psi(t)=x_{0}$, which implies that the insurer enters the market with initial capital $x_{0}$, commencing all business activities ((re)insurance and investment) precisely at time $t=0$. Therefore, the initial condition for $M_{1}^{u}(t)$ is $M_{1}^{u}(0)=\frac{x_{0}(1-e^{-\alpha h})}{\alpha}$. Furthermore, we adopt the following notations for convenience. For $t\in[-h,0]$, denote $M_{1}^{u}(t)=m_{10}$ and $M_{2}^{u}(t)=m_{20}$. For $\tau\in[0,T]$ and $t\in[\tau-h,\tau]$, denote $X^{u}(t)=x$, $M_{1}^{u}(t)=m_{1}$ and $M_{2}^{u}(t)=m_{2}$.

### v2

@@ Abstract
This paper considers a newly delayed reinsurance and investment optimization problem incorporating random risk aversion, in which an insurer pursues maximization of the expected certainty equivalent of her/his terminal wealth and the cumulative delayed information of the wealth over a period. Specially, the insurer’s surplus dynamics are approximated using a drifted Brownian motion, while the financial market is described by the constant elasticity of variance (CEV) model. Moreover, the performance-linked capital flow feature is incorporated and the wealth process is formulated via a stochastic delay differential equation (SDDE). By adopting a game-theoretic approach, a verification theorem with rigorous proofs is established to capture the equilibrium reinsurance and investment strategy along with the equilibrium value function. Furthermore, analytical or semi-analytical equilibrium reinsurance and investment strategies, together with their equilibrium value functions, are obtained under the CEV model for the exponential utility and derived under the Black-Scholes model for both exponential and power utilities. Finally, several numerical experiments are conducted to analyze the behavioral characteristics of the freshly-derived equilibrium reinsurance and investment strategy.
@@ 1 Introduction
Risk management has always been a crucial subject for insurers. Reinsurance serves as an efficient mean for insurers to manage claim-related risk exposure and safeguard themselves against catastrophic loss scenarios. Within a reinsurance contract, an insurer transfers a specified portion of claim losses to a reinsurer and remits the reinsurance premium for the risk transfer. Conversely, the reinsurer has the obligation to reimburse a portion of the claims that the insurer has incurred. Generally, the magnitude of reinsurance premium rises in line with the proportion of ceded claims, which compels the insurer to strike a balance between the reinsurance premium and ceded claims in order to attain the optimally achieved state of the reinsurance arrangement. Moreover, nowadays, insurers are actively engaged in investment pursuits within financial markets. Thus, it is also highly relevant to formulate suitable investment strategies to boost returns to the fullest extent or/and cut down risks to the minimum when dealing with different investment tools.
A substantial number of studies have been conducted on formulating optimal reinsurance and investment strategies for insurers. In these studies, many scholars have dedicated to expanding the previous works to include a variety of optimality criteria and distinct foundational models for the insurer’s surplus dynamics and risky asset evolution. For example, Browne [13] studied the optimal portfolio allocation problem for an insurer, taking into account the criteria of utility maximization and minimization of ruin probability separately; Højgaard and Taksar [22] aimed at enabling an insurer to pursue maximization of the return function for the proportional reinsurance problem; Yang and Zhang [34] adopted a jump-diffusion model to capture an insurer’s surplus process and investigated the optimal investment issue. Many more studies and outcomes associated with the reinsurance and investment problems are available in the literature (see, for instance, Bai and Guo [4], Elliott and Siu [18], Li et al. [27], Wang et al. [32] and the references therein).
@@
One typical feature of the above researches is that these researches concentrated on the controlled state system in the absence of delay considerations. Specially, it was commonly supposed that the system’s future state has no connection with past states and is governed exclusively by the present situation. Nevertheless, in real-world systems, insurers often take past information into account when formulating their reinsurance and investment strategies, as past information can help insurers understand the cyclical patterns of markets and enhance the robustness of decision-making. This feature is frequently termed as the memory feature or the delay feature. Chang et al. [15] used the HJB equation to study a consumption and investment problem featuring bounded memory, where the return of the risky asset was associated with the historic performance. Then, Shen and Zeng [30] pioneered the incorporation of bounded wealth memory into a reinsurance and investment problem subject to the mean-variance criterion, deriving the pre-committed strategy through the stochastic maximum principle. It was pointed out in [15, 30] that since the delayed optimal control problems are usually infinite-dimensional, a significant challenge emerges, namely, the non-existence of the explicit solutions. Therefore, some specific constraints on model parameters were given in [15, 30] to render the original problems finite-dimensional and solvable. Further research works and results related to delayed optimal control problems (including delayed optimal reinsurance/investment problems) are well-documented in the literature (see, for example, A et al. [1], A and Shao [2], A et al. [3], Bai et al. [5], Elsanosi et al. [19], Federico [20], Meng et al. [28], Yan and Wong [33], Yuan et al. [38] and the references therein).
@@
It is worthy of note that the majority of aforementioned studies assumed that investors/insurers had a constant risk aversion coefficient. In reality, investors/insurers cannot easily determine their exact level of risk aversion. This has motivated extensive research on risk aversion estimation through diverse methodologies such as employing data from the market on labor income [16], using data pertaining to implied and realized volatilities from the market [12] and examining survey forms of representative samples [14]. Recently, Desmettre and Steffensen [17] proposed that it is more reasonable to characterize the risk aversion coefficient of investors with a random variable and they formulated a new investment optimization problem in the Black-Scholes financial market by employing expected certainty equivalent, which is time inconsistent. They also provided a strict verification theorem to capture the equilibrium investment strategy, deriving (semi-)analytical strategies for both power and exponential utility specifications. More recently, based on the research of [17], Kang et al. [25] considered an optimal reinsurance and investment problem incorporating random risk aversion in the financial market characterized by the Heston stochastic volatility model. They also presented the pseudo HJB equation to characterize the equilibrium reinsurance and investment strategy, from which the semi-analytical equilibrium strategy was obtained under the exponential utility.
@@
To the best knowledge of the authors, there is currently no literature that adopts the method of expected certainty equivalent to study the delayed optimal reinsurance and investment problem under random risk aversion. Thus, this paper systematically addresses this gap by applying expected certainty equivalent to consider a new delayed optimal reinsurance and investment problem incorporating random risk aversion. In particular, under the diffusion approximation modeling of the insurer’s surplus process, the insurer can cede risk exposure to a reinsurer by employing proportional reinsurance strategy or seek growth through new business expansion and can also optimize asset allocations in a financial market following the CEV model. Furthermore, it is hypothesized that capital inflows and outflows affect the insurer’s present wealth, with the associated wealth dynamics formulated via an SDDE. Then, an approach based on expected certainty equivalent is employed to formulate the general delayed optimal reinsurance and investment problem considering random risk aversion that can be characterized as a combination of nonlinear functions of expectations.
@@
The above newly established problem exhibits time inconsistency because Bellman’s dynamic programming principle is inapplicable. In order to obtain the (semi-)analytical equilibrium reinsurance and investment strategy for such a problem, we identify the following challenges. Firstly, due to the introduction of delay, the methodologies in [17, 25] for providing the verification theorem are no longer directly applicable. In this paper, we draw on [19] to derive the Itô formula under delay scenario, and then adopt the methodologies from [9, 11, 17, 25], integrating them with the aforementioned Itô formula to provide a verification theorem in delay context. Secondly, as discussed earlier, the inclusion of delay leads to the problem being infinite-dimensional. To overcome this difficulty, we follow [15, 30] to impose specific assumptions on model parameters to transform the problem into a finite-dimensional framework, which facilitates our derivation of (semi-)analytical equilibrium reinsurance and investment strategies under the exponential utility and the power utility. Finally, we find that only a small number of studies, such as [5, 15], have proven the admissibility of their delayed optimal (or equilibrium) strategies and the conditions in their verification theorem. However, these proofs have certain limitations and cannot be directly applied to this paper. For example, substituting equations (56) and (57) into equation (12) in [15], we discover that the proof process of Lemma 2.2 is no longer applicable and thus $c^{*}(t)\geq 0$ in the proof of Theorem 4.2 cannot be easy to check. What is more, Bai et al. [5] did not consider the verification of the integrability of the value function (see the assumption $(A1)$ in this paper). To address this issue, we utilize the properties of solutions to stochastic differential equations (SDEs) and the assumption of model parameters to complete the corresponding proofs. After overcoming these challenges, we would like to point out that the newly derived equilibrium reinsurance and investment strategies extend some results in [1, 2, 17, 25] and also arrive at several novel and interesting conclusions. For instance, the expressions of the equilibrium reinsurance strategy and the equilibrium investment strategy under the power utility function in the Black-Scholes financial market demonstrate that even under the independence assumption between financial and insurance markets, both strategies are unexpectedly influenced by the model parameters from the reinsurance and financial market setups.
@@
The primary contributions of this paper can be outlined from four aspects: (i) a newly delayed optimal reinsurance and investment model with random risk aversion is established, which fully generalizes the model in [17] and partially generalizes the model in [25], while also enriching the research on time-inconsistent problems under delay scenarios; (ii) a complete verification theorem is provided to capture the equilibrium reinsurance and investment strategy and the associated equilibrium value function; (iii) analytical or semi-analytical equilibrium reinsurance and investment strategies, together with their equilibrium value functions, are obtained under the CEV model for the exponential utility and derived under the Black-Scholes model for both exponential and power utilities, followed by the proofs of their admissibility and the validity of the verification theorem’s conditions; (iv) computational experiments quantify parameter sensitivity in the equilibrium reinsurance and investment strategy, which yields novel and intriguing findings.
@@
The subsequent sections of this paper are structured as outlined below. Section 2 details the model framework and sets forth the fundamental assumptions. Section 3 poses the delayed optimal reinsurance and investment problem under random risk aversion and gives a verification theorem capturing the equilibrium reinsurance and investment strategy along with the equilibrium value function. Section 4 derives the (semi-)closed-form solutions to equilibrium reinsurance and investment strategies under the CEV model for the exponential utility. Section 5 obtains (semi-)closed-form equilibrium reinsurance and investment strategies under the Black-Scholes model for exponential and power utilities separately. Section 6 presents numerical simulations to quantify how model parameters influence the equilibrium reinsurance and investment strategy. Section 7 offers a comprehensive summary of this study and Appendices include the detailed proofs.
@@ 2.2 Financial market
where $r>0$ is a constant denoting the risk-free interest rate. In addition, the price process $S_{1}(t)$ of the risky asset is governed by the CEV model, which satisfies
$\mathrm{d}S_{1}(t)=\mu S_{1}(t)\mathrm{d}t+\sigma S^{\delta+1}_{1}(t)\mathrm{d}W_{2}(t),\quad S_{1}(0)=\bar{s}_{1}>0,$
where $\mu$ and $\sigma S^{\delta}_{1}(t)$ characterize respectively the expected return rate and instantaneous volatility of $S_{1}(t)$ with $\mu>r$ and $\sigma>0$, $\delta>0$ denotes the elasticity parameter, and $W_{2}:=\{W_{2}(t)\}_{t\in[0,T]}$ is a standard $\mathbb{P}$-Brownian motion independent of $W_{1}$. We note that the CEV model degenerates into geometric Brownian motion when $\delta=0$.
@@ 2.3 Wealth process with delay
$\mathrm{d}X^{u}(t)$ $=\frac{(1-\pi(t))X^{u}(t)}{S_{0}(t)}\mathrm{d}S_{0}(t)+\frac{\pi(t)X^{u}(t)}{S_{1}(t)}\mathrm{d}S_{1}(t)+\mathrm{d}R(t)$
$=[X^{u}(t)\left(r+\pi(t)(\mu-r)\right)+a\eta+a\eta_{2}q(t)]\mathrm{d}t+bq(t)\mathrm{d}W_{1}(t)+\pi(t)X^{u}(t)\sigma S^{\delta}_{1}(t)\mathrm{d}W_{2}(t),$
where the wealth value at time 0 is $X^{u}(0)=x_{0}>0$.
@@
$\mathrm{d}X^{u}(t)=$ $[X^{u}(t)\left(r+\pi(t)(\mu-r)\right)+a\eta+a\eta_{2}q(t)]\mathrm{d}t+bq(t)\mathrm{d}W_{1}(t)+\pi(t)X^{u}(t)\sigma S^{\delta}_{1}(t)\mathrm{d}W_{2}(t)$
@@
$+bq(t)\mathrm{d}W_{1}(t)+\pi(t)X^{u}(t)\sigma S^{\delta}_{1}(t)\mathrm{d}W_{2}(t),$ (2.3)
@@
We suppose that the initial condition for $X^{u}(t)$ is its historical performance from $\tau-h$ to $\tau$ formulated as $X^{u}(t)=\psi(t-\tau)$ for $t\in[\tau-h,\tau]$, where $\psi>0$ is a continuous function on [h,0][-h,0]. If $\tau=0$, then we suppose that the insurer’s initial wealth is constant, namely, for any $t\in[-h,0]$, $X^{u}(t)=\psi(t)=x_{0}$, which implies that the insurer enters the market with initial capital $x_{0}$, commencing all business activities ((re)insurance and investment) precisely at time $t=0$. Therefore, the initial condition for $M_{1}^{u}(t)$ is $M_{1}^{u}(0)=\frac{x_{0}(1-e^{-\alpha h})}{\alpha}$. Furthermore, we adopt the following notations for convenience. For $t\in[-h,0]$, denote $M_{1}^{u}(t)=m_{10}$ and $M_{2}^{u}(t)=m_{20}$. For $\tau\in[0,T]$ and $t\in[\tau-h,\tau]$, denote $S_{1}(t)=s_{1}$, $X^{u}(t)=x$, $M_{1}^{u}(t)=m_{1}$ and $M_{2}^{u}(t)=m_{2}$.